SafeュDriving Taxi Service charges $5 per pickup and $0.25 per mile. We値l Get You There Cabs charges $1
per pickup and $0.50 per mile. At how many miles will the two companies charge you the same amount?

Answer by rfer(16322)

5+0.25x=1+0.50x
0.25x-0.50x=1-5
-0.25x=-4
x=-4/-0.25
x=16 mi